# Exploiting Inter-Session Information with Frequency-enhanced Dual-Path Networks for Sequential Recommendation

### v1

@@ Abstract
Sequential recommendation (SR) aims to predict a user’s next item preference by modeling historical interaction sequences. Recent advances often integrate frequency-domain modules to compensate for self-attention’s low-pass nature by restoring the high-frequency signals critical for personalized recommendations. Nevertheless, existing frequency-aware solutions process each session in isolation and optimize exclusively with time-domain objectives. Consequently, they overlook cross-session spectral dependencies and fail to enforce alignment between predicted and actual spectral signatures, leaving valuable frequency information under-exploited. To this end, we propose FreqRec, a Frequency-Enhanced Dual-Path Network for sequential Recommendation that jointly captures inter-session and intra-session behaviors via a learnable Frequency-domain Multi-layer Perceptrons. Moreover, FreqRec is optimized under a composite objective that combines cross entropy with a frequency-domain consistency loss, explicitly aligning predicted and true spectral signatures. Extensive experiments on three benchmarks show that FreqRec surpasses strong baselines and remains robust under data sparsity and noisy-log conditions. Our code is available at: https://github.com/AONE-NLP/FreqRec.
@@ Introduction
Figure 1: The distribution of Pearson correlation coefficient between sessions sharing at least one item on Sports & Outdoors and Beauty. Calculation details provide in Appendix.
Recent work has explored two complementary strategies to address these limitations in frequency domain. One line of work (zhou2022filter; du2023contrastive; liu2023distribution; kim2025diff) applies explicit spectral filtering to suppress spurious signals and highlight essential sequence features. Another line (du2023frequency; shin2024attentive) integrates the Fourier transform modules alongside self-attention to recover periodic patterns and complementary high-frequency information. Despite of their state-of-the-art performance, frequency-based methods also introduce new challenges while answering the aforementioned questions:
@@
Sequential recommendation (SR) aims to learn from users’ historical interaction sequences to deliver personalized recommendations. Current mainstream SR models predominantly adopt deep learning frameworks: Convolutional Neural Networks (CNNs) (tang2018personalized; niu2024diffusion) and Recurrent Neural Networks (RNNs) (li2018learning; liu2018stamp) were initially widely employed. For instance, Caser (tang2018personalized) pioneered the treatment of items in recommendation sequences, utilizing convolutional operations to capture item-item interactions. GRU4Rec (Hidasi2015SessionbasedRW) successfully implemented a GRU network to construct a SR framework for modeling long-term dependencies between users-items. Transformers (vaswani2017attention), owing to their exceptional long-term contextual modeling capabilities and parallel computation efficiency, have gained widespread adoption in SR. Models such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec) use self-attention to adaptively weight historical interactions, thereby more effectively capturing shifts in user intent. However, the global receptive field of self-attention acts as a low-pass filter, smoothing out abrupt changes of interests in user behavior (du2023frequency).
To recover these crucial spectral cues, a growing body of work integrates frequency-domain processing within SR models. One line of research designs explicit spectral filters to attenuate noise and amplify informative frequency bands. For instance, Zhou et al. (zhou2022filter) propose a learned band-pass filter that adapts its frequency response to sequence statistics, suppressing spurious interactions while retaining periodic patterns. SLIME4Rec (du2023contrastive) employs contrastive learning on spectral representations to improve robustness against noisy logs, while DIFF (kim2025diff) leverages DFT-based filtering, combined with side information, to remove short-term fluctuations and enhance recommendation accuracy. A second strand embeds Fourier transform operations directly into the network architecture to recover high-frequency dynamics that self-attention may smooth over. FEARec (du2023frequency) and BSARec (shin2024attentive) utilized DFT to capture high-frequency information, counteracting the inherent low-pass filtering properties of self-attention mechanisms. Although these methods achieve strong empirical gains, they process each session in isolation and optimize solely with time-domain losses. These gaps motivate our development of a dual-path architecture that simultaneously models inter-session and intra-session spectral dependencies while incorporating a frequency-domain consistency objective.
@@
Instead of hand-crafting a low-pass or band-pass filter, we propose a Frequency-Domain MLP (FreqMLP) that learns which frequencies to amplify or suppress. As illustrate in Figure 3, FreqMLP jointly processes the real and imaginary channels, enabling cross-component interactions:
@@
Local Spectral Refiner (LSR). Complementing the GSA stage, the LSR capture each user’s fine-grained and often transient interests by applying the same frequency filtering along the temporal dimension $L$, given the sequence embedding of users $\mathcal{E}_{\mathrm{Intra}}\in\mathbb{R}^{B\times L\times D}$:
@@
  • Parallel Fusion processes inter- and intra-session streams independently and combines them with a weighted sum, which preserves the integrity of the raw sequence features and allows each module to specialize without mutual interference:
@@
We first apply the DFT $\mathcal{F}$ to transform both the predicted output $P$ and target values $T$ into the frequency domain, respectively obtaining their real and imaginary components. Subsequently, we compute the frequency distance function $\mathcal{F_{D}}(P,T)$ (as detailed in Appendix) to evaluate the discrepancies between their real components and between their imaginary components in the frequency domain. This process is formally defined as:
@@
We conducted a comprehensive evaluation on three classic SR datasets  [^1] https://cseweb.ucsd.edu/ jmcauley/datasets/amazon_v2/, Amazon Beauty, Sports & Outdoors, and Toys & Games. We borrowed the data preprocessing approach from previous work (shin2024attentive; kim2025diff), and details about the description of the datasets can be found in Appendix.
@@
Implementation Details.
All experiments are implemented in PyTorch and executed on a server equipped with an NVIDIA RTX 4090 GPU (24 GB). To ensure comparability, we adopted the hyper-parameter settings of prior work (shin2024attentive; kim2025diff). Specifically, we set the hidden dimension $D$ to 64 and the maximum sequence length $L$ to 50. Optimization was performed using Adam with learning rates in $\{\text{5e-4},\text{1e-4},\text{1e-3}\}$, while a 10-round early-stopping criterion is deployed. The resulting optimal hyper-parameters are reported in the Appendix.
@@
    A comparative evaluation against Large Language Model (LLM) and Generative models, along with detailed computational efficiency analysis, is provided in Appendix.
@@
Table 3: Impact of the frequency-domain consistency loss ($\mathcal{L_{F}}$). Red superscripts indicate the improvement rates. More results can be found in the Appendix.
Effectiveness of Frequency Domain Loss
In this section, we assess the effectiveness and flexibility of our Fourier loss $\mathcal{L_{F}}$ by integrating it into four representative sequential recommendation baselines on the Beauty and Toys & Games datasets. To ensure experimental fairness, we adhere strictly to each model’s original implementation and hyper-parameter settings. As shown in Table 3, our proposed Fourier loss consistently improves baseline performance in a true plug‐and‐play fashion, recording average improvements over both datasets in terms of HR@10 and NDCG@10 ranging from 3.70% to 19.45%. It is noticeable that models which already exploit frequency domain features (FMLPRec and BSARec) derive the greatest benefit. FMLPRec’s HR@10 climbs by 15.21% (from 0.0559 to 0.0644) and NDCG@10 by 19.93% (from 0.0291 to 0.0349), while BSARec secures a 46.45% gain in NDCG@10 (from 0.0437 to 0.0639). These considerable improvements underscore that explicitly penalizing frequency discrepancies provides a complementary signal to cross-entropy loss, guiding DFT-based models toward more discriminative feature learning. We detail the choices of distance calculation protocols for frequency loss in Appendix.
@@
Table 4: Performance under noisy scenario in terms of HR@10 and NDCG@10 across the three target domains. Extended results are provided in the Appendix.
Figure 4: Hyperparametric sensitivity experiments on the Beauty and Toys & Games datasets.
@@
We benchmark FreqRec under the noise-resilience ability following the protocol of previous work (liu2025preference). Instead of the standard supervised paradigm, we train the model on the aggregate interaction logs from three Amazon categories, Automotive, CDs & Vinyl, and Grocery & Gourmet Food, and then evaluate it on each category independently, with no fine-tuning on the target domain. This evaluation setting requires the model to distinguish meaningful behavioral signals from heterogeneous and domain-specific noise in the training data. Statistical descriptions of the datasets are provided in Appendix.
@@
Table 5: Performance under sparse interaction degrees.
Performance w.r.t Data Sparsity
@@ Conclusion
In this paper, we present FreqRec, a Frequency-Enhanced Dual-Path Network for sequential recommendation that jointly models inter-session and intra-session dynamics via a learnable, complex-valued Fourier transform. By coupling a Global Spectral Aggregator for cohort-level rhythms with a Local Spectral Refiner for user-specific signals, FreqRec effectively suppresses noise and restores high-frequency behavioral patterns that standard attention mechanisms smooth out. We further introduce a frequency-domain consistency loss, which explicitly aligns the model’s predicted spectral coefficients with ground-truth signatures, thereby bridging the gap left by conventional time-domain objectives. Extensive experiments on three public benchmarks show that FreqRec delivers state-of-the-art performance, while maintains robustness under data sparsity and noisy logs scenarios.
@@ Acknowledgements
We would like to thank the anonymous reviewers for their valuable disucssion and constructive feedback. This work is supported by National Natural Science Foundation of China (U22B2061), National Key R&D Program of China (2022YFB4300603), and Natural Science Foundation of Sichuan, China (2024NSFSC0496).

### v3

@@ Abstract
Sequential recommendation (SR) aims to predict a user’s next item preference by modeling historical interaction sequences. Recent advances often integrate frequency-domain modules to compensate for self-attention’s low-pass nature by restoring the high-frequency signals critical for personalized recommendations. Nevertheless, existing frequency-aware solutions process each session in isolation and optimize exclusively with time-domain objectives. Consequently, they overlook cross-session spectral dependencies and fail to enforce alignment between predicted and actual spectral signatures, leaving valuable frequency information under-exploited. To this end, we propose FreqRec, a Frequency-Enhanced Dual-Path Network for sequential Recommendation that jointly captures inter-session and intra-session behaviors via a learnable Frequency-domain Multi-layer Perceptron. Moreover, FreqRec is optimized under a composite objective that combines cross entropy with a frequency-domain consistency loss, explicitly aligning predicted and true spectral signatures. Extensive experiments on three benchmarks show that FreqRec surpasses strong baselines and remains robust under data sparsity and noisy-log conditions. Our code is available at: https://github.com/AONE-NLP/FreqRec.
@@ Introduction
Figure 1: The distribution of Pearson correlation coefficient between sessions sharing at least one item on Sports & Outdoors and Beauty datasets, respectively.
Recent work has explored two complementary strategies to address these limitations in the frequency domain. One line of work (zhou2022filter; du2023contrastive; liu2023distribution; kim2025diff) applies explicit spectral filtering to suppress spurious signals and highlight essential sequence features. Another line (du2023frequency; shin2024attentive) integrates the Fourier transform modules alongside self-attention to recover periodic patterns and complementary high-frequency information. Despite their state-of-the-art performance, frequency-based methods introduce new challenges while answering the aforementioned questions:
@@
Sequential recommendation (SR) aims to learn from users’ historical interaction sequences to deliver personalized recommendations (gan2025pareto). Mainstream SR models predominantly adopt deep learning frameworks: Convolutional Neural Networks (CNNs) (tang2018personalized; niu2024diffusion) and Recurrent Neural Networks (RNNs) (li2018learning; liu2018stamp) were initially widely employed. For instance, Caser (tang2018personalized) pioneered the treatment of items in recommendation sequences, utilizing convolutional operations to capture item-item interactions. GRU4Rec (Hidasi2015SessionbasedRW) successfully implemented a GRU network to construct a SR framework for modeling long-term dependencies between users-items. Transformers (vaswani2017attention), owing to their exceptional long-term contextual modeling capabilities and parallel computation efficiency, have gained widespread adoption in SR. Models such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec) use self-attention to adaptively weight historical interactions, thereby more effectively capturing shifts in user intent. However, the global receptive field of self-attention acts as a low-pass filter, smoothing out abrupt changes in interests in user behavior (du2023frequency).
To recover these crucial spectral cues, a growing body of work integrates frequency-domain processing within SR models. One line of research designs explicit spectral filters to attenuate noise and amplify informative frequency bands. For instance, Zhou et al. (zhou2022filter) propose a learned band-pass filter that adapts its frequency response to sequence statistics, suppressing spurious interactions while retaining periodic patterns. SLIME4Rec (du2023contrastive) employs contrastive learning on spectral representations to improve robustness against noisy logs, while DIFF (kim2025diff) leverages DFT-based filtering, combined with side information, to remove short-term fluctuations and enhance recommendation accuracy. A second strand embeds Fourier transform operations directly into the network architecture to recover high-frequency dynamics that self-attention may smooth over. FEARec (du2023frequency) and BSARec (shin2024attentive) utilize DFT to capture high-frequency information, counteracting the inherent low-pass filtering properties of self-attention mechanisms. Although these methods achieve strong empirical gains, they process each session in isolation and optimize solely with time-domain losses. These gaps motivate our development of a dual-path architecture that simultaneously models inter-session and intra-session spectral dependencies while incorporating a frequency-domain consistency objective.
@@
Instead of hand-crafting a low-pass or band-pass filter, we propose a Frequency-Domain MLP (FreqMLP) that learns which frequencies to amplify or suppress. As illustrated in Figure 3, FreqMLP jointly processes the real and imaginary channels, enabling cross-component interactions:
@@
Local Spectral Refiner (LSR). Complementing the GSA stage, the LSR captures each user’s fine-grained and often transient interests by applying the same frequency filtering along the temporal dimension $L$, given the sequence embedding of users $\mathcal{E}_{\mathrm{Intra}}\in\mathbb{R}^{B\times L\times D}$:
@@
  • Parallel Fusion processes inter- and intra-session streams independently and combines their respective outputs with a weighted sum, which preserves the integrity of the raw sequence features and allows each module to specialize without mutual interference:
@@
We transform the prediction $P$ and target $T$ into the frequency domain. We define $\mathcal{L}_{\mathrm{L1}}(P,T)=\frac{1}{N}\sum_{i=1}^{N}|P_{i}-T_{i}|$, $\mathcal{L}_{\mathrm{L2}}(P,T)=\frac{1}{N}\sum_{i=1}^{N}(P_{i}-T_{i})^{2}$, and $\mathcal{L}_{\mathrm{mix}}=\frac{1}{2}(\mathcal{L}_{\mathrm{L1}}+\mathcal{L}_{\mathrm{L2}})$. The distance function $\mathcal{F_{D}}$ is selected from:
$\mathcal{F_{D}}(P,T)\in\{\mathcal{L}_{\mathrm{L1}}(P,T),\mathcal{L}_{\mathrm{L2}}(P,T),\mathcal{L}_{\mathrm{mix}}(P,T)\}$ (18)
where $P,T\in\mathbb{R}^{B\times L\times D}$. The final frequency-domain loss $\mathcal{L_{F}}$ separately computes the distance between the real and imaginary parts of $P$ and $T$ after the DFT $\mathcal{F}$:
@@
We conducted a comprehensive evaluation on three widely adopted SR benchmarks [^1] https://cseweb.ucsd.edu/ jmcauley/datasets/amazon_v2/: Beauty, Sports & Outdoors, and Toys & Games. These datasets, derived from the Amazon e-commerce platform, contain product reviews from 1996 to 2014 and are widely used for the SR task. We adopted the data preprocessing approach from previous work (shin2024attentive; kim2025diff).
@@
Implementation Details
Experiments were conducted on an NVIDIA RTX 4090 (24 GB) using PyTorch. We employed the Adam optimizer with 10-round early stopping, and a learning rate grid-searched from $\{\text{5e-4},\text{1e-4},\text{1e-3}\}$. For fair comparison, we followed prior work (shin2024attentive; kim2025diff) setting $D=64$ and $L=50$. For FreqRec, we grid-searched $\gamma,\alpha\in\{0.1,0.3,0.5,0.7,0.9\}$, $\beta\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$, batch size $B\in\{32,64,128,256,512\}$, and $\mathcal{F_{D}}\in\{\mathcal{L}_{\mathrm{L1}},\mathcal{L}_{\mathrm{L2}},\mathcal{L}_{\mathrm{mix}}\}$.
@@
Table 3: Impact of the frequency-domain consistency loss ($\mathcal{L_{F}}$). Red superscripts indicate the improvement rates.
@@
Table 4: Performance under noisy scenario in terms of HR@10 and NDCG@10 across the three target domains.
Effectiveness of Frequency Domain Loss
In this section, we assess the effectiveness and flexibility of our Fourier loss $\mathcal{L_{F}}$ by integrating it into four representative sequential recommendation baselines on the Beauty and Toys & Games datasets. To ensure experimental fairness, we adhere strictly to each model’s original implementation and hyper-parameter settings. As shown in Table 3, our proposed Fourier loss consistently improves baseline performance in a true plug-and-play fashion, recording average improvements over both datasets in terms of HR@10 and NDCG@10 ranging from 3.70% to 19.45%. It is noteworthy that models which already exploit frequency domain features (FMLPRec and BSARec) derive the greatest benefit. FMLPRec’s HR@10 climbs by 15.21% (from 0.0559 to 0.0644) and NDCG@10 by 19.93% (from 0.0291 to 0.0349), while BSARec secures a 46.45% gain in NDCG@10 (from 0.0437 to 0.0639). These considerable improvements underscore that explicitly penalizing frequency discrepancies provides a complementary signal to cross-entropy loss, guiding DFT-based models toward more discriminative feature learning.
Figure 4: Sensitivity test of hyper-parameters on the Beauty and Toys & Games datasets.
@@
We benchmark FreqRec under the noise-resilience ability following the protocol of previous work (liu2025preference). Instead of the standard supervised paradigm, we train the model on the aggregate interaction logs from three Amazon categories, Automotive, CDs & Vinyl, and Grocery & Gourmet Food, and then evaluate it on each category independently, with no fine-tuning on the target domain. This evaluation setting requires the model to distinguish meaningful behavioral signals from heterogeneous and domain-specific noise in the training data.
@@
Table 5: Performance of FreqRec against baselines under different sparse interaction degrees.
@@ Conclusion
In this paper, we present FreqRec, a Frequency-Enhanced Dual-Path Network for sequential recommendation that jointly models inter-session and intra-session dynamics via a learnable, complex-valued Fourier transform. By coupling a Global Spectral Aggregator for cohort-level rhythms with a Local Spectral Refiner for user-specific signals, FreqRec effectively suppresses noise and restores high-frequency behavioral patterns that standard attention mechanisms smooth out. We further introduce a frequency-domain consistency loss, which explicitly aligns the model’s predicted spectral coefficients with ground-truth signatures, thereby bridging the gap left by conventional time-domain objectives. Extensive experiments on three public benchmarks show that FreqRec delivers state-of-the-art performance, while maintaining robustness under data sparsity and noisy logs scenarios.
@@ Acknowledgements
We would like to thank the anonymous reviewers for their valuable discussion and constructive feedback. This work is supported by National Natural Science Foundation of China (U22B2061), National Key R&D Program of China (2022YFB4300603), and Natural Science Foundation of Sichuan, China (2024NSFSC0496).